OpenStudy (anonymous):

Simplify. Please be sure to show all of your work. 8-5/8+1/2*6/9

jhonyy9 (jhonyy9):

8 - (5/8) + (1/2) x (6/9) = (64-5)/8 + (6/18) = 59/8 + 1/6 = ((3 x 59) + 4)/24 = = (177 +4)/24 = 181/24